What volume will 12 g of fluorine gas occupy at stp

1
Expert's answer
2021-11-15T10:56:28-0500

At STP p= 1atm , T = 273.15, R= 8.314,

"PV = nRT"

n= mass/ rfm

n = 12/19= 0.6315

V= (0.6315× 273.15× 8.314)/1

V= 1434L